OpenStudy (anonymous):

Please help me solve the following ODE (explicit soln): dy/dx = (y-1)^2 , y(0) = 1 I have worked through it and my answer is y = -1/x but my book is showing the answer as y = 1 ):

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=(y-1)^{2} , y(0) = 1\]

OpenStudy (anonymous):

Applying separating variables method

OpenStudy (anonymous):

\[\frac{ dy }{ (y-1)^2 }=dx\]

OpenStudy (anonymous):

\[\int\limits (\frac{ dy }{ (y-1)^2 }=dx)\]

OpenStudy (anonymous):

\[\int\limits\limits \frac{1}{(y-1)^2} = \int\limits\limits dx\]

OpenStudy (anonymous):

\[-\frac{1}{y-1}=x+c\]

OpenStudy (anonymous):

ok we need an alternative approach because as we can see this can not be solve by implicit ODE

OpenStudy (anonymous):

so let's try the explicit ODE...

OpenStudy (anonymous):

What is the explicit ODE?

OpenStudy (anonymous):

I think the explicit form is \[y=-\frac{ 1 }{ x }-1+c\] which doesn't help me with the initial condition either. So I tried only simplifying to the form \[-1=x(y-1)+c\] to get c = -1, which if I plug into my explicit form, makes the answer \[y=-\frac{ 1 }{ x }\]

OpenStudy (anonymous):

That still doesn't agree with my book's answer of y=1

OpenStudy (anonymous):

yes because solving this explicitly requires recalling the exponential function in order to get the right answer

OpenStudy (anonymous):

if the explicit method does not give a 1 then we will consider that an error in the book

OpenStudy (anonymous):

by the y(0)=1 which means x=0

OpenStudy (anonymous):

right, so using the explicit method with that initial condition makes it undefined...right?

OpenStudy (anonymous):

no using implicit method makes it undefined

OpenStudy (anonymous):

Then what is the explicit method?

OpenStudy (anonymous):

ok i got y = 1

OpenStudy (anonymous):

sorry i was trying this on a piece of paper

OpenStudy (anonymous):

here is how you do it

OpenStudy (anonymous):

The suspense is killing me! (:

OpenStudy (anonymous):

sorry i was on the phone

OpenStudy (anonymous):

ok we integrate the 1/(y-1)^2 by expanding the value 1/(y^2-2y +1), then we integrate RHS we get x+c

OpenStudy (anonymous):

we will end up with y=quberoot(3x-1+c)+1, which yields to the fact that y=1 for x=0 and y(0)=1

OpenStudy (anonymous):

i need to restart

OpenStudy (anonymous):

the pc brb

OpenStudy (anonymous):

LHS still integrates to -1/(y-1)